OpenStudy (anonymous):

Can I do this? a² = b² + c² 13² = 12 ² + c² c² = 13² - 12² c = sqtr(13 - 12) c = sqrt(1) c = 1

OpenStudy (zzr0ck3r):

sqrt(1) = +-1

jimthompson5910 (jim_thompson5910):

are you trying to find the missing leg given a leg of 12 and a hypotenuse of 13?

OpenStudy (zzr0ck3r):

and you cant do what you did with the sqrt

OpenStudy (zzr0ck3r):

c² = 13² - 12² c = sqtr(13^2 - 12^2)

jimthompson5910 (jim_thompson5910):

if so, then a^2 + b^2 = c^2 12^2 + b^2 = 13^2 solve for b

OpenStudy (zzr0ck3r):

\[c^{2} = a^{2} + b^{2}\] then \[c = \sqrt(a^{2}+b^{2}\]

OpenStudy (anonymous):

Ok. I got. Thank you all.